Blood contains positive and negative ions, which experience a magnetic force in a magnetic field as they move. An Electromagnetic Flowmeter takes advantage of this phenomenon. A heart surgeon monitors the flow rate of blood through an artery. Electrodes make contact with the outer surface of the blood vessel to measure an induced voltage. For a magnetic field strength of 0.037 T, a potential difference of 160 µV appears between the electrodes for a blood vessel of inside diameter 2.9 mm.

a) How fast is the blood flowing in the artery?



Artery To N S Electrodes Bloed llow
Answer #1

Potential difference electrodes will be:

emf = B*v*L*sinA

here, B = magnetic field = 0.037 T

v = speed of blood in artery = ??

L = width of artery = diameter = 2.9 mm = 2.9*10^-3 m

A = angle between velocity vector and magnetic field vector = 90 deg

given, emf = 160 uV = 160*10^-6 V

then, v = emf/(B*L*sinA)

v = (160*10^-6)/(0.037*2.9*10^-3*sin(90 deg))

v = 1.49 m/s

"Let me know if you have any query."